I do not understand how to write the expressions for finding y or how to find the amounts of ingredients in this problem:
Suppose that a nurse needs 14 cc (cubic centimeters) of an 18.5% solution (that is, a solution that is 18.5% ingredient) of a certain medication, which must be obtained by mixing x cc of a 20% solution and y cc of a 9% solution. Write expressions involving x for the following:
y= ____
the amount of ingredient in the x cc of 20% solution_____
the amount of ingredient in the 9% solution ______
the total amount of ingredient in the mixture _______

Answer by mananth(16946)

component percent ---------------- quantity
x 9 ---------------- x cc
y 20 ---------------- 14 - x cc
Mixture 18.50 ---------------- 14

9 x + 20 (14 -x) = 18.50 * 14.00

9 x + 280 - 20 x= 259.00
9 x - 20 x= 259 - 280
-11 x = -21
/ -11
x = 1.90 cc Solution I (x)
12.09 cc Solution II (y)